OpenStudy (anonymous):

OK I NEED HELP LIKE ALOT! How far is Town Center from Maple Station? (Points : 2) 4 2 5.83 5.57

OpenStudy (kymber):

Needs a diagram

OpenStudy (anonymous):

IS THERE ANY WAY I CAN SEND YOU THE PIC OF IT?

OpenStudy (kymber):

Yes.. press "attach file"

OpenStudy (anonymous):

did u get it?

OpenStudy (anonymous):

you need to use the distance formula... Maple Station is 3 units to the left and 5 units higher than the Town Center. So the distance formula is: Distance between 2 points = square root ( a^2 + b^2) where "a" and "b" are horizontal and vertical distances... in this problem, "a" and "b" would be 3 and 5, so you use those in the distance formula. \[distance = \sqrt{ a^{2} + b^{2} }\]

OpenStudy (anonymous):

so its 8?

OpenStudy (anonymous):

no, you have to square "a" and "b", then add, then take the square root.

OpenStudy (anonymous):



OpenStudy (anonymous):

34?

OpenStudy (anonymous):

3^2 = 9 5^2 = 25 3^2 + 5^2 = 9 + 25 = 34 Now take the square root of 34...